OpenStudy (paradox2018):

Evaluate the function f(x)=2x + 2, when x=1

OpenStudy (jtug6):

f(x) = y = 2x+2, and we need to know what y is when x =1. So, everywhere you see x, just plug 1 in it's place. y = 2(1) + 2. What is 2(1) + 2?

OpenStudy (paradox2018):

4

OpenStudy (jtug6):

yup. so, f(x) = 2x + 2 at x = 1 is 4. So, f(x) = y = 4

OpenStudy (paradox2018):

Thanks, Can you help with one other one? Equations are hard for me lol

OpenStudy (jtug6):

Yeah def. Ask away

OpenStudy (paradox2018):

Evaluate the function g(t)=5/2t - 3/2, when t = 3

OpenStudy (joshtray):

Just as jtug6 has said, wherever you see t, just put in 3 You will have\[\frac{ 5 }{ 2\times3 }-\frac{ 3 }{ 2 }\]

OpenStudy (jtug6):

Yup! Exactly. Sorry for late response was helping someone else

OpenStudy (joshtray):

2 x 3 is 6 So then we can proceed by finding the L.C.M of the two denominators

OpenStudy (paradox2018):

so, 2?

OpenStudy (paradox2018):

6 is one of my possible answers

OpenStudy (joshtray):

The L.C.M. of 6 and 2 is what?

OpenStudy (paradox2018):

Oh, Duh, 6

OpenStudy (joshtray):

So we will then have\[\frac{ 5-9 }{ 6 }\]which is\[\frac{-4}{6}\] which is equivalent to\[-\frac{2}{3}\]

OpenStudy (paradox2018):

Thanks!